OpenStudy (anonymous):

find the volume of the solid obtained by rotating the region bounded by the given curves about the specified line. sketch the region, the solid, and a typical disk or washer. y=x^3, y=0, x=1; about x=2

OpenStudy (anonymous):

draw the graph of X^3 is the first step

OpenStudy (anonymous):

as well as y=0, X=1, and X=2

OpenStudy (anonymous):

yes i have the graph, im just not sure how to set up the integral

OpenStudy (anonymous):

the volume of 1 washer =\[\pi(R-r)^{2}\Delta X\]

OpenStudy (anonymous):

what is R? R= x^3+1

OpenStudy (anonymous):

and what would r be?

OpenStudy (anonymous):

the limit is from 0 to 1. and we rotate is about x=2.  that is the small r .

OpenStudy (anonymous):

to get r we will have to subtract 2-1=1 r=1